# How Quantum Complexity Helps Classical Complexity: A Survey

### 16 Dec 2023 v1

@@ 6. Conclusion
In conclusion, the revelations of quantum mechanics have illuminated a distinct realm, governed by its own principles. Quantum computing stands out as a promising avenue to transcend the limitations of classical computing, unlocking solutions to previously insurmountable problems. However, challenges such as uncontrolled errors pose obstacles to its full potential.
@@
Various research groups have undertaken endeavors to tackle classical computational problems using quantum devices. In our exploration, we have compiled papers that delve into this approach and pinpointed key open problems demanding attention.
@@
    Quantum Supremacy: It is a common question that quantum computers can outperform classical computers, as discussed in problems 1950, and 104. What is quantum supremacy? Is it limited to a particular type of problem or universal? How can we demonstrate that quantum computers can solve problems that are practically infeasible tasks for classical computers?
@@
    Quantum Error Correction: Managing and controlling qubits is difficult and the system can be easily disturbed by noise and errors during computation. Detection and correction of these errors is a challenging task, but it is necessary to make quantum computers more reliable and scalable. Problems 612263797, and 105 are example of this type of challenge.
@@
    Quantum Algorithms: Using quantum hardware instead of classical ones requires new algorithms that are compatible with these devices. To harness the supremacy of quantum computing, we should design algorithms that are more efficient than classical ones. Some examples of such algorithms are provided in problems: 615161723384967, and 103.
@@
    Hybrid Quantum-Classical Computing: Combining classical and quantum computers might be a powerful approach to solving hard problems. This has usually been achieved by assigning different tasks to each type of computer respectively. Problems 14 and 39 are examples of papers that used hybrid computation in their research.
@@
    Quantum Simulation: Quantum systems are complex and have several variables which make them difficult to simulate using classical computers. However, simulating a quantum system using another quantum system could be easier and more efficient. This could open up new possibilities for materials science, chemistry, and more. problems 32 and 33 are examples of problems related to this topic.
@@
    Quantum Cryptography: Quantum computation could impact cryptography significantly due to the unique properties of quantum mechanics like entanglement. Also, it’s unknown How quantum computing can enhance security protocols, such as quantum key distribution and secure communication. Problems 4445545585, and 86 are about quantum cryptography.
@@
    Scalability: Quantum technologies could be revolutionary technology if they could be scalable. Nowadays controlling quantum systems and reducing noises are not fully available so the number of qubits in quantum computers is small. As it mention in problems 25814, and 35, building larger and more stable quantum system is desirable.
@@
    Quantum Software and Programming Languages: As said before, quantum computers work with distinct algorithms. As a result, the embracement of quantum computers requires developing user-friendly programming languages and tools. It is desirable to have a quantum processor that can run any program to find the solution, although it seems impossible to have such an exact programable processor. Problem 9 mentions this issue.
@@
    Quantum Annealing: Quantum annealing is a technique for computation that finds the state with minimum energy in the system. This strategy is different from gate-based computation and is hard to compare. Problem 9 discuss about this case.
@@
    Quantum Complexity Theory: It seems that the definition of complexity of quantum computation is different from the classical one. Quantum computational complexity classes are unique and their relation with classical complexity classes is not known. Problems: 43596365666869707375, and 77 are about this subject.
Note that these open problems not only expand our understanding of quantum computation but also hold the key to unlocking its full potential. The intersection of quantum computing and classical computation offers a vast landscape for exploration. Researchers can develop new solutions and broaden the horizons of what is possible.
@@ 7.3. Authors’ contributions
The idea was for the first author. All of the other authors are sorted based on their contribution.

### 10 Jan 2024 v2

@@ 6. Conclusion
In conclusion, the revelations of quantum mechanics have illuminated a distinct realm, governed by its principles. Quantum computing stands out as a promising avenue to transcend the limitations of classical computing, unlocking solutions to previously insurmountable problems. However, challenges such as uncontrolled errors pose obstacles to its full potential.
@@
Various research groups have undertaken endeavors to tackle classical computational problems using quantum devices. In our exploration, we have compiled papers that delve into this approach and pinpoint key open problems demanding attention.
@@
    Quantum Supremacy: The question of whether quantum computers can surpass classical computers (quantum supremacy) is a common one, and is explored in problems 1950, and 104. The concept of quantum supremacy is multifaceted and raises several questions: What precisely is quantum supremacy? Is it restricted to certain types of problems or is it all-encompassing? How can we demonstrate that quantum computers can solve problems that are practically infeasible tasks for classical computers?
    To demonstrate that quantum computers can solve problems that are practically impossible for classical computers to handle, one must establish a method of proof. This can be accomplished through the use of formal proof or through empirical evidence gathered from experimental data. As quantum computers continue to develop and improve, the demonstration of quantum supremacy will likely become increasingly feasible.
@@
    Quantum Error Correction: The management and control of qubits presents a formidable challenge in the context of quantum computing, given the system’s susceptibility to noise and errors during computation. The detection and correction of these errors is an essential task to enhance the reliability and scalability of such computing systems. Collectively, problems 612263797, and 105 represent exemplars of the challenges that arise in this arena, and underscore the importance of developing effective strategies for addressing them.
@@
    Quantum Algorithms: To employ quantum hardware instead of classical alternatives, it is necessary to develop algorithms that are congruent with these devices. To fully harness the supremacy of quantum computing, it is imperative that we design algorithms that are more efficient than their classical counterparts. Examples of such algorithms are presented in problems:  615161723384967, and 103.
@@
    Hybrid Quantum-Classical Computing: The integration of classical and quantum computers can be a potent approach to solving intricate problems. This is typically achieved by allocating specific tasks to each type of computer. Two exemplary research papers that utilized hybrid computation in their work are Problems 14 and 39. In light of the complex nature of contemporary problems, the amalgamation of classical and quantum computing has emerged as a promising avenue for addressing hitherto intractable computational challenges. The judicious allocation of tasks between classical and quantum computers can lead to significant improvements in computational efficiency and accuracy.
@@
    Quantum Simulation: Quantum systems are characterized by their complexity and multi-dimensional nature, rendering their simulation on classical computers challenging. Nonetheless, the simulation of a quantum system utilizing another quantum system may offer a more efficient alternative. This avenue of approach could potentially unlock novel possibilities in the fields of materials science, as well as chemistry. Two illustrative examples of problems germane to this subject matter are problems 32 and 33.
@@
    Quantum Cryptography: Quantum computation has the potential to significantly impact cryptography due to the unique properties of quantum mechanics, such as entanglement. However, the extent to which quantum computing can enhance security protocols, such as quantum key distribution and secure communication, remains unclear. Notwithstanding, numerous problems in quantum cryptography have been identified, as enumerated by problems 4445545585, and 86.
@@
    Scalability: Quantum technologies have the potential to be revolutionary, but their scalability remains a challenge. Currently, controlling quantum systems and reducing noise is not fully feasible, resulting in a limited number of qubits in quantum computers. As problems 25814, and 35 highlight, building larger and more stable quantum systems is essential for their advancement.
@@
    Quantum Software and Programming Languages: As previously stated, quantum computers operate using distinctive algorithms. Consequently, the integration of quantum computers into practical applications necessitates the development of user-friendly programming languages and tools. It is highly desirable to have a quantum processor capable of executing any program to obtain the solution. However, the development of such a precisely programmable processor appears to be an insurmountable challenge. This issue is discussed in depth in Problem 9.
@@
    Quantum Annealing: Quantum annealing is a computational technique that aims to find the state with the minimum energy in a given system. Unlike gate-based computation, quantum annealing is distinct and cannot be readily compared with it. More details are discussed in Problem 9.
@@
    Quantum Complexity Theory: Quantum computational complexity classes have a unique definition that differs from the classical one. The relationship between these classes and classical complexity classes is unknown. Problems 43 to 77 are related to this topic. Problems: 43596365666869707375, and 77 are about this subject.
It is imperative to acknowledge that the unresolved problems in quantum computation not only serve to advance our understanding of the field but also play a pivotal role in unlocking its full potential. The intersection of quantum computing and classical computation presents a vast and uncharted research landscape. Researchers have the opportunity to generate innovative solutions and expand the scope of what is achievable. By addressing these open questions, we can pave the way for further advancements in the field of quantum computation, ultimately leading to a greater understanding of complex computational problems.
@@ 7.3. Authors’ contributions
The idea was for the first author. All of the other authors are sorted based on their contributions.